Question 1


Find the derivative of

y = ln 3x5.

 

a.    ln 3x5

b.    15x4

c.    5x−1

d. 1/

3x5

 


Question  2

Find the derivative of the function


a.    x2/ x + x/ x

b. x2 + x /x

c. 2x + 1

d.    3x4 + 2x3

 


Question  3

Differentiate the function

f (x) = 3x3e2x.


a. 5x4e3x

b. 3x2e2x(x − 3)

c. 3e3x + 9x4x5

d. 3x2e2x(2x + 3)


Question  4

The total revenue function (TR), in terms of P and the company's quantity, if a commodity sells for R199.00  are?

a. TR = 6 000 − 30P ; Q = 30.

b. TR = 6 000P − 30P 2; Q = 30.

c. TR = P (6 000 − 30P ); Q = 24.

d.  TR = 6 000 − 30P ; Q = 24.

 



1
Expert's answer
2022-01-10T14:38:19-0500

1.

"y' =\frac{15x^4}{3x^5}=5x^{-1}"


Answer: c. 5x−1


2.

"y=x^2+x"

"y'=2x+1"


Answer: c. 2x + 1


3.

"f' (x) =9x^2e^{2x}+ 6x^3e^{2x}=3x^2e^{2x}(2x + 3)"


Answer: d. 3x2e2x(2x + 3)


4.

The total revenue function (TR), in terms of P and the company's quantity, if a commodity sells for R199.00 are?

"TR = Q P = 6 000 P - 30 P^2"

"Q = 6 000 - 30 P = 6 000 - 30 \cdot199 = 30"


Answer: b. TR = 6 000P − 30P 2; Q = 30